# Fluid flow inside slit-shaped nanopores: the role of surface morphology at the molecular scale

### v1

@@ Abstract
Non-equilibrium molecular dynamics (NEMD) simulations of fluid flow have highlighted the peculiarities of nanoscale flows compared to classical fluid mechanics. In particular, boundary conditions can deviate from the no-slip behavior at macroscopic scales due to various factors. In this context, we investigate the influence of surface morphology in slit-shaped nanopores on the fluid flow. We demonstrate that the surface morphology effectively controls the slip length, which approaches zero when the molecular structures of the pore wall and the fluid are matched. Using boundary-driven, energy-conserving NEMD simulations with a pump-like driving mechanism, we examine two types of pore walls—mimicking a crystalline and an amorphous material—that exhibit markedly different surface resistances to flow. The resulting flow velocity profiles are consistent with Hagen–Poiseuille theory for incompressible, Newtonian fluids when adjusted for surface slip. For the two pores, we observe partial slip and no-slip behavior, respectively, which correlate with fluid layering and depletion near the surfaces. However, the confinement of the fluid gives rise to an effective viscosity that varies substantially with the pore width. Analysis of the hydrodynamic permeability shows that the simulated flows are in the Darcy regime. Additionally, the thermal isolation of the flow causes a linear increase in fluid temperature along the flow, which we relate to strong viscous dissipation and heat convection, utilizing conservation laws of fluid mechanics. Our findings underscore the need for molecular-scale modeling to accurately capture the fluid dynamics near boundaries and in nanoporous materials, where macroscopic models may not be applicable.
@@ I Introduction
In the past 60 years, molecular dynamics (MD) simulations have developed into an indispensable tool in many areas of research, complementing experimental investigations and theoretical predictions; examples range from drug development [1] and cell biology [2] to materials research [3, 4], chemical process engineering [5, 6], and geology [7]. Nowadays, it is used routinely to study the physical and chemical properties of complex materials and macromolecules as well as kinetic processes therein [8, 9, 10, 11]. This provides detailed insights into phase diagrams, structural properties, transformation rates, and transport coefficients, including dynamic shear viscosity, thermal and electrical conductivity, and diffusivity. The pioneering works [12, 13] employed classical particle models with simplistic interaction potentials to test fundamental predictions of statistical mechanics and kinetic theory. In this respect, the results of MD simulations have set and are still setting corner stones in our understanding of material responses, which were supported by and are still challenging the theoretical developments, a prime example being the long-time tails in the autocorrelation of particles in liquids [14, 15, 16, 17] and arrested fluids [18, 19, 20]. Moreover, the development of advanced, realistic force fields [21, 22, 23, 24] has turned MD simulations into a prime method for quantitative predictions of material properties, and these developments have rapidly gained new momentum with the advent of machine-learned interaction potentials that are based on first-principles calculations [25, 26, 27].
@@
Whereas the investigated system sizes were originally limited to a few hundred atoms and just several thousand integration time steps, simulations of some million atoms over tens of millions of time steps can be performed routinely on today’s high-performance computing installations [28, 29, 30, 31], also owing to the massive parallelism provided by general-purpose accelerator hardware (“graphics processors”). This has brought small-wave number and low-frequency phenomena into the reach of MD simulations [32, 33, 16, 34, 28], thereby providing access to hydrodynamic scales with molecular resolution. It allows one to study set-ups typical of fluid mechanics in great detail such that also the boundary layers of a fluid flow can be resolved down to the scale of atomic interactions. This is particularly relevant for the development of realistic continuum models of pores via up-scaling approaches, which need information about the molecular behaviour at the nanoscale [35, 36]; there, hydrodynamic permeability models are prone to break down [37].
@@
Transport phenomena in confined geometries, particularly at the nanoscale, showcase a continuously growing number of technological applications and relevance in natural processes [38, 39, 40]. The behavior of fluids in confined spaces is studied with the aim to design materials for filtration [41, 42], energy applications [37, 42, 43, 44], water desalination [45], chip-level cooling [46], and DNA sequencing [47], to name only a few. A critical factor in determining fluid behavior within a porous medium is the nature of the interactions between the fluid and the pore surfaces [48, 49]. Hydrophobicity, the tendency of a surface to repel water or similar polar fluids, modulates the fluid flow and distribution within the pore space; in particular, it can lead to pronounced slippage of the flow at the fluid–surface interface. Such slip, however, depends in a subtly way on factors such as surface roughness, wettability, and the presence of gaseous layers [50]. Surface roughness can change the fluid behavior significantly, which is known as the Lotus effect [49]: experiments evidence that specific surface coatings—such as polymers or nanotube structures, referred to as “nanoforests”—can create superhydrophobic surfaces with extremely low friction to flow [51]. The effect has also been demonstrated in MD simulations of fluid flow near nano-structured surfaces [52]. On the other hand, for hydrophilic surfaces with a similar topography, one may expect that trapping of fluids within surface asperities or pockets creates additional friction and reduces slippage.
@@
For atomically smooth, crystalline surfaces, it was shown that hydro- and solvophobicity facilitate fluid slip [53], which was linked to the formation of distinct interfacial layers, where fluid density and structure differ significantly from those in the bulk [54]. However, despite a general trend that hydrophobic surfaces show a reduced fluid–surface interaction and enhance fluid slip, hydrophobicity alone does not imply a large slip length and vice versa [55]. Moreover, subtle changes in the fluid–surface interaction can drive significant structural and density changes near the surface without modifying the wetting properties of the surface. There is no simple correlation between the interfacial fluid structure and the hydrophobicity of the surface and we refer to the partially controversial discussion in, e.g., Refs. 56, 57, 58. Also atomic-level details like surface composition, molecular orientation, and electronic interactions can profoundly impact the slip behavior, even on structurally similar surfaces, which was demonstrated in ab initio simulations of water near graphene surfaces [59, 60].
@@
Fluid flow through nanoporous media using non-equilibrium MD (NEMD) simulations has been investigated with low intensity since the 1990’s (e.g., Ref. 61), but has seen a surge in research activity during the past decade. The focus is typically on the permeability of ordered porous media such as MOFs [62, 63] and linear or slit-shaped pores [54, 64, 65, 66, 67, 68], but also through disordered organic materials [69, 70, 37]. The majority of these studies addresses the transport of molecules in the gas phase or of low-density fluids, while NEMD simulations for dense liquids face particular technical challenges. One approach to steady fluid flow is to prescribe the chemical potential of the inlet and outlet by means of grand-canonical (GC) reservoirs (dual control volume set-up) [71, 69, 70]. However, the Monte Carlo part of this approach would become inefficient for dense liquids due to the low acceptance rate of trial insertions of particles. One solution to this problem is to control the pressure instead of the chemical potential [72], and a recent, reservoir-based alternative to GC-NEMD simulations is offered by spatially adaptive resolution simulations (AdResS) [73, 74, 75]. Another line of approach to steady flow is driving the liquid in a boundary region of a periodic set-up by applying an external force to maintain a pressure difference [67, 68] or a concentration gradient [63]. More generally, custom flow fields can be generated by extending the force field into the whole pore space [76]; in the context of incompressible flows, a constant force was used to mimic a constant pressure gradient [54, 65]. Below, we describe boundary-driven NEMD simulations of steady flow implement via a pump-like driving mechanism, which is applied to the particles in a small region at the entrance of the pore, while preserving the Hamiltonian character of the fluid dynamics in the remaining pore volume.
@@
In this work, we investigate how surface morphology, or molecular-scale surface topography, affects fluid flow near the surface. In particular, how does it modify the effective boundary conditions, the local structure of the fluid, and the overall permeability? We propose and characterize a nanopore model with discrete walls composed of obstacles arranged to match the molecular structure of the fluid. We show that such a configuration allows efficient control of surface slip, which is almost completely eliminated. More specifically, we simulate the steady flow of a simple fluid inside a slit-shaped nanopore as a minimal model of a porous medium; this choice enables comparison with previous studies and has obvious extensions to other pore cross sections. We relate the molecular-level observations of fluid flow to the basic laws of fluid mechanics and test predictions of Hagen-Poiseuille theory for incompressible, Newtonian fluids. Lastly, we discuss the issue of heat transport in isolated fluid flow under conditions of strong viscous dissipation, as encountered in fast-flow scenarios.
II Fluid mechanical considerations
II.1 Hagen–Poiseuille flow in a slit-shaped conduit
@@
We consider the laminar flow of a fluid through a slit-like conduit of width $D$ and square-shaped surfaces, each of macroscopic area $L^{2}$. The application of a pressure difference $\Delta P=P_{\text{in}}-P_{\text{out}}$ at the inlet and outlet of the conduit leads to a Hagen–Poiseuille flow, which we shall briefly re-derive here. To this end, we choose a coordinate frame such that the $x$-axis points parallel to the wall surfaces, in the direction of the flow, and the $z$-axis is perpendicular to the surfaces, which are located at $z=\pm D/2$ so that $z=0$ denotes the plane in the center of the flow. By translational symmetry, the problem does not depend on the $y$-coordinate of the position.
The steady flow of a uniform, incompressible, Newtonian fluid is governed by the stationary Navier–Stokes equation [77, 78]:
$m\varrho\left({\boldsymbol{v}}\cdot\nabla\right){\boldsymbol{v}}=-\nabla P+\eta\nabla^{2}{\boldsymbol{v}},$ (1)
with $m\varrho(x,z)$ being the fluid density field, ${\boldsymbol{v}}(x,z)$ the velocity field, $P(x,z)$ the pressure field, and $\eta$ the dynamic shear viscosity. At macroscopic scales, no-slip boundary conditions hold at the surfaces of the conduit:
${\boldsymbol{v}}(x,\pm D/2)=0.$ (2)

### v2

@@ Abstract
Non-equilibrium molecular dynamics (NEMD) simulations of fluid flow have highlighted the peculiarities of nanoscale flows compared to classical fluid mechanics; in particular, boundary conditions can deviate from the no-slip behavior at macroscopic scales. For fluid flow in slit-shaped nanopores, we demonstrate that surface morphology provides an efficient control on the slip length, which approaches zero when matching the molecular structures of the pore wall and the fluid. Using boundary-driven, energy-conserving NEMD simulations with a pump-like driving mechanism, we examine two types of pore walls—mimicking a crystalline and an amorphous material—that exhibit markedly different surface resistances to flow. The resulting flow velocity profiles are consistent with Poiseuille theory for incompressible, Newtonian fluids when adjusted for surface slip. For the two pores, we observe partial slip and no-slip behavior, respectively. The hydrodynamic permeability corroborates that the simulated flows are in the Darcy regime. However, the confinement of the fluid gives rise to an effective viscosity below its bulk value; wide pores exhibit a crossover between boundary and bulk-like flows. Additionally, the thermal isolation of the flow causes a linear increase in fluid temperature along the flow, which we relate to strong viscous dissipation and heat convection, utilizing conservation laws of fluid mechanics. Noting that the investigated fluid model does not form droplets, our findings challenge the universality of previously reported correlations between slippage, solvophobicity, and a depletion zone. Furthermore, they underscore the need for molecular-scale modeling to accurately capture the fluid dynamics near boundaries and in nanoporous materials, where macroscopic models may not be applicable.
@@ I Introduction
In the past 60 years, molecular dynamics (MD) simulations have developed into an indispensable tool in many areas of research, complementing experimental investigations and theoretical predictions; examples range from drug development [1] and cell biology [2] to materials research [3, 4], chemical process engineering [5, 6], and geology [7]. Nowadays, they are used routinely to study the physical and chemical properties of complex materials and macromolecules as well as kinetic processes therein [8, 9, 10, 11]. This provides detailed insights into phase diagrams, structural properties, transformation rates, and transport coefficients, including dynamic shear viscosity, thermal and electrical conductivity, and diffusivity. The pioneering works [12, 13] employed classical particle models with simplistic interaction potentials to test fundamental predictions of statistical mechanics and kinetic theory. In this respect, the results of MD simulations have set and are still setting corner stones in our understanding of material responses, which were supported by and are still challenging the theoretical developments, a prime example being the long-time tails in the autocorrelation of particles in liquids [14, 15, 16, 17] and arrested fluids [18, 19, 20]. Moreover, the development of advanced, realistic force fields [21, 22, 23, 24] has turned MD simulations into a prime method for quantitative predictions of material properties, and these developments have rapidly gained new momentum with the advent of machine-learned interaction potentials that are based on first-principles calculations [25, 26, 27].
@@
Whereas the investigated system sizes were originally limited to a few hundred atoms and just several thousand integration time steps, simulations of some million atoms over tens of millions of time steps can be performed routinely on today’s high-performance computing installations [28, 29, 30, 31], also owing to the massive parallelism provided by general-purpose accelerator hardware (“graphics processors”). This has brought small-wave number and low-frequency phenomena within the reach of MD simulations [32, 33, 16, 34, 28], thereby giving access to hydrodynamic scales with molecular resolution. It allows one to study set-ups typical of fluid mechanics in great detail such that also the boundary layers of a fluid flow can be resolved down to the scale of atomic interactions. This is particularly relevant for the development of realistic continuum models of pores via up-scaling approaches, which need information about the molecular behavior at the nanoscale [35, 36]; there, hydrodynamic permeability models are prone to break down [37].
@@
Transport phenomena in confined geometries, particularly at the nanoscale, showcase a continuously growing number of technological applications and relevance in natural processes [38, 39, 40]. The behavior of fluids in confined spaces is studied with the aim to design materials for filtration [41, 42], energy applications [37, 42, 43, 44], water desalination [45], chip-level cooling [46], and DNA sequencing [47], to name only a few. A critical factor in determining fluid behavior within a porous medium is the nature of the interactions between the fluid and the pore surfaces [48, 49]. Hydrophobicity, the tendency of a surface to repel water or similar polar fluids, modulates the fluid flow and distribution within the pore space; in particular, it can lead to pronounced slippage of the flow at the fluid–wall interface. Such slip, however, depends in a subtly way on factors such as surface roughness, wettability, and the presence of gaseous layers [50]. Surface roughness can change the fluid behavior significantly, which is known as the Lotus effect [49]: experiments evidence that specific surface coatings—such as polymers or nanotube structures, referred to as “nanoforests”—can create superhydrophobic surfaces with extremely low friction to flow [51]. The effect has also been demonstrated in MD simulations of fluid flow near nano-structured surfaces [52]. On the other hand, for hydrophilic surfaces with a similar topography, one may expect that trapping of fluids within surface asperities or pockets creates additional friction and reduces slippage.
@@
For atomically smooth, crystalline surfaces, it was shown that hydro- and solvophobicity facilitate fluid slip [53, 54], which was linked to the formation of distinct interfacial layers, where fluid density and structure differ significantly from those in the bulk [55]. However, despite a general trend that hydrophobic surfaces show a reduced fluid–surface interaction and enhance fluid slip, hydrophobicity alone does not imply a large slip length and vice versa [56]. Moreover, subtle changes in the fluid–surface interaction can drive significant structural and density changes near the surface without modifying the wetting properties of the surface. There is no simple correlation between the interfacial fluid structure and the hydrophobicity of the surface and we refer to the partially controversial discussion in, e.g., Refs. 57, 58, 59. Also atomic-level details like surface composition, molecular orientation, and electronic interactions can profoundly impact the slip behavior, even on structurally similar surfaces, which was demonstrated in ab initio simulations of water near graphene surfaces [60, 61].
@@
Fluid flow through nanoporous media using non-equilibrium MD (NEMD) simulations has been investigated with low intensity since the 1990’s (e.g., Ref. 62), but has seen a surge in research activity during the past decade. The focus is typically on the permeability of ordered porous media such as MOFs [63, 64] and linear or slit-shaped pores [55, 65, 66, 67, 68, 69], but also through disordered organic materials [70, 71, 37]. The majority of these studies addresses the transport of molecules in the gas phase or of low-density fluids, while NEMD simulations for dense liquids face particular technical challenges. One approach to steady fluid flow is to prescribe the chemical potential of the inlet and outlet by means of grand-canonical (GC) reservoirs (dual control volume set-up) [72, 70, 71]. However, the Monte Carlo part of this approach would become inefficient for dense liquids due to the low acceptance rate of trial insertions of particles. One solution to this problem is to control the pressure instead of the chemical potential [73], and a recent, reservoir-based alternative to GC-NEMD simulations is offered by spatially adaptive resolution simulations (AdResS) [74, 75, 76]. Another line of approach to steady flow is driving the liquid in a boundary region of a periodic set-up by applying an external force to maintain a pressure difference [68, 69] or a concentration gradient [64]. More generally, custom flow fields can be generated by extending the force field into the whole pore space [77]; in the context of incompressible flows, a constant force was used to mimic a constant pressure gradient [55, 66]. Below, we describe boundary-driven NEMD simulations of steady flow implemented via a pump-like driving mechanism, which is applied to the particles in a small region at the entrance of the pore, while preserving the Hamiltonian character of the fluid dynamics in the remaining pore volume.
@@
In this work, we investigate how surface morphology, or molecular-scale surface topography, affects fluid flow near the surface. In particular, how does it modify the effective boundary conditions, the local structure of the fluid, and the overall permeability? We propose and characterize a nanopore model with discrete walls composed of obstacles arranged to match the molecular structure of the fluid. We show that such a configuration allows efficient control of surface slip, which is almost completely eliminated. More specifically, we simulate the steady flow of a simple fluid inside a slit-shaped nanopore as a minimal model of a porous medium; this choice enables comparison with previous studies and has obvious extensions to other pore cross sections. We relate the molecular-level observations of fluid flow to the basic laws of fluid mechanics and test predictions of Poiseuille theory for incompressible, Newtonian fluids. Lastly, we discuss the issue of heat transport in isolated fluid flow under conditions of strong viscous dissipation, as encountered in fast-flow scenarios [78].
II Simulation model
Figure 1: Top: sketch of the geometry of the simulated slit-shaped pore space. The cuboid is filled with fluid, which is driven at the pore entrance (yellow region) in the direction of the arrows, indicating the flow. The flow is delimited by the pore walls at the top and bottom faces of the cuboid, and periodic boundary conditions are applied at the other faces.  Bottom: Snapshots of the particle configurations in NEMD simulations for two different obstacle arrangements (red particles), which realize walls with a crystalline (fcc) and amorphous structure, respectively. The pore is filled with fluid particles (blue). For better visibility, a smaller pore space ($L=50$) than the one used in the actual simulations is shown; the inset is a close-up of the fluid–wall interface.
We consider the fluid flow confined to a nanoscale conduit of slit geometry, which is created by two parallel, square-shaped walls of nominal distance $D$ and each of surface area $L\times L$, Fig. 1. The volume accessible to the fluid has the shape of a cuboid, of which two opposite open faces are referred to as the inlet and outlet of the conduit; periodic boundary conditions are used at the remaining two faces. The application of a pressure difference $\Delta P=P_{\text{in}}-P_{\text{out}}$ between the inlet and outlet leads to plane Poiseuille flow; equivalently, one may impose a mean flow velocity on the fluid and observe the resulting pressure drop.
@@ II Simulation model
We choose a coordinate frame such that the $x$-axis points parallel to the wall surfaces, in the direction of the flow, and the $z$-axis is normal to the surfaces, which are located at $z=\pm D/2$ so that $z=0$ denotes the plane in the center of the flow, Fig. 1a). By translational symmetry, the problem does not depend on the $y$-coordinate of the position.
II.1 Fluid model
which is derived from the Lennard-Jones pair potential, $U_{\text{LJ}}(r)=4\varepsilon\bigl{[}(\sigma/r)^{12}-(\sigma/r)^{6}\bigr{]}.$ The parameters $\varepsilon$ and $\sigma$ serve as units of energy and length, respectively, and the unit of time is chosen to be $\tau=\sqrt{m\sigma^{2}/\varepsilon}$ in terms of the particle mass $m$. Placing the cutoff distance $r_{c}$ at the minimum of the potential, $r_{c}=2^{1/6}\sigma$, the interaction contains only the repulsive forces that prevent particle overlap without considering attractive dispersion forces. A smooth truncation is used to ensure good energy conservation during long MD simulations in the microcanonical (NVE) ensemble, implemented via multiplication of the energy-shifted potential with the algebraic interpolation function [79, 80, 81, 33] $g(x)=x^{4}/\bigl{(}1+x^{4}\bigr{)}$ for $x\leqslant 0$ and $g(x)=0$ otherwise. Thereby, $U_{\text{WCA}}(r)$ is rendered $C^{2}$-continuous also at the cutoff. We choose $h=0.005\sigma$ for the truncation range, which is small enough to entail only minor changes of the fluid properties due to the truncation. Due to the purely repulsive interaction, the fluid does not exhibit distinguishable liquid and vapor phases and, therefore, it does not form stable droplets in equilibrium.
[…]
temperature $T$=\>=\>= 1.0 1.0 3.0 $\varepsilon/k_{\text{B}}$
mass density $m\varrho$=\>=\>= 0.7 0.9 0.55 $m\sigma^{-3}$
pressure $P$=\>=\>= 4.277 10.130 5.360 $\varepsilon\sigma^{-3}$
internal energy $e$=\>=\>= $1.462(2)$ $1.334(5)$ $4.419(3)$ $\varepsilon/m$
isochoric specific heat $c_{V}$=\>=\>= $2.034(1)$ $2.527(5)$ $1.791(1)$ $k_{\text{B}}/m$
compressibility $\chi_{T}$=\>=\>= $0.04256(1)$ $0.00991(2)$ $0.0600(1)$ $\sigma^{3}/\varepsilon$
speed of sound $c_{s}$=\>=\>= $5.79(1)$ $10.58(1)$ $5.50(1)$ $\sigma/\tau$
shear viscosity $\eta_{b}$=\>=\>= $1.423(2)$ $3.494(6)$ $4.17(2)$ $\sigma^{3}\tau/\varepsilon$
thermal conductivity $\kappa$=\>=\>= $9.38(3)$ $13.51(5)$ $9.15(4)$ $k_{\text{B}}/(\sigma\tau)$
Table 1: Thermodynamic observables, material constants, and transport coefficients of the purely repulsive, WCA fluid, at three thermodynamic state points $(\varrho,T)$ representative of the flowing fluid; the quiescent fluid corresponds to $T=1.0\varepsilon/k_{\text{B}}$ and $\varrho=0.7\sigma^{-3}$. The values were obtained with equilibrium MD simulations following standard procedures, the speed of sounds was calculated from $c_{s}=(\varrho\chi_{T})^{-1/2}$. Numbers in parentheses give the uncertainty in the last digit.
II.2 Pore morphologies
The pore walls are implemented as arrangements of obstacle particles, which are fixed in space and which exert forces on the fluid particles according to the same WCA potential, Eq. 1, that is used for the interaction between fluid particles. The use of such discrete walls provides us with flexibility regarding the surface morphology. This allows us to study the role of morphology in the fluid flow by comparing simulations with two types of pore surfaces, which we refer to as crystalline and as amorphous walls, Fig. 1.
For crystalline walls, the pore particles are placed on a face-centered cubic (fcc) lattice with a lattice constant of $a_{\text{lat}}=1.15\sigma$ such that the lattice planes are located at $z=\pm[D/2+(n-1/2)a_{\text{lat}}/2]$ for integer $n\geqslant 1$. Technically, the dimension of the simulation box is extended along the $z$-axis by adding $15\sigma$ in total to accommodate the obstacle particles. The lattice constant is chosen small enough to prevent infiltration of fluid particles into the wall structure, it implies nearest-neighbor distances of $a_{\text{lat}}/\sqrt{2}\approx 0.82\sigma$. The dense arrangement of pore particles results in a highly ordered, crystalline structure, exposing a comparably smooth surface to the fluid.
For amorphous walls, the pore particles are distributed in a disordered fashion in such a way that the fluid structure is matched; a similar approach was used earlier to study confinement effects in supercooled liquids [82, 83]. Such wall structures are generated by filling the whole simulation volume, $(D+15\sigma)\times L\times L$, with a WCA fluid at the desired fluid density and temperature and equilibrating its structure in a short NVT simulation run. Afterwards, the particles found outside the (nominal) volume of the conduit are identified as obstacles and they delimit the pore space, while those inside the pore region represent the fluid; obstacle particles do not participate in the further integration of the equations of motion. The procedure leads to comparably rough, irregular wall surfaces in contact with the fluid.
Due to the presence of density fluctuations in the obstacle configuration and the overall lower obstacle density, a small number of fluid particles can penetrate deep into the pore walls and get “lost”. As an additional measure to ensure particle conservation in the pore volume, we placed two flat, impenetrable walls alongside the obstacles at positions $z_{\text{flat}}^{\pm}=\pm(D/2+\sigma)$, inside of the pore walls. For the interaction energy of a fluid particle at position $z$ with each wall, we use a purely repulsive surface potential:
in terms of the distances $\hat{z}=z-z_{\text{flat}}^{+}$ and $z=z-z_{\text{flat}}^{-}$. The surface potential is smoothly truncated, analogously to Eq. 1, using again the cutoff distance $r_{c}=2^{1/6}\sigma$ with $h=0.005\sigma$. We emphasize that these additional, flat potentials are positioned at a distance of $\sigma$ well behind the pore surfaces and that the particles of the moving fluid are essentially not interacting with them.
III Fluid mechanical considerations
III.1 Poiseuille flow in a slit-shaped conduit
In fluid mechanics, the laminar flow through a slit-like conduit is referred to as plane Poiseuille flow [84, 85], which develops in response to an applied pressure difference $\Delta P$ along the conduit. The latter is of length $L$ and width $D$ and the extension in the remaining direction (along the $y$-axis) is taken to be macroscopically large (alternatively, periodic boundary conditions can be applied, as we have done in the NEMD simulations). Here, we shall briefly recall the main results of the theory with an emphasis on the underlying assumptions.